# Revisiting Semi-Supervised Learning in the Era of Foundation Models

### v2

@@ 1 Introduction
Recently, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains gui2024survey . For example, in computer vision, many recent algorithms are built upon vision foundation models (VFMs) such as CLIP radford2021learningtransferablevisualmodels and DINOv2 oquab2024dinov2learningrobustvisual . These models were pre-trained on massive datasets—millions, if not billions, of data points. As a result, they have demonstrated remarkable generalizability across a wide range of tasks, often requiring only slight fine-tuning or, in some cases, functioning effectively as a frozen backbone zhang2024vision .
@@ 2 Related Works
Transfer Learning & Self-Supervised Learning. Transfer learning zhuang2020comprehensivesurveytransferlearning —and specifically PEFT (or PETL) for efficiently adapting foundation models han2024parameterefficientfinetuninglargemodels —has long leveraged pretrained models to boost downstream tasks. In our approach, we fine-tune VFMs via PEFT on labeled data, then use the adapted models to generate pseudo-labels for unlabeled examples, thus bridging transfer learning with semi-supervised methods. A related paradigm, self-supervised learning, also exploits unlabeled data and often serves as a baseline in SSL comparisons chen2020bigselfsupervisedmodelsstrong ; assran2022maskedsiamesenetworkslabelefficient , but it assumes abundant unlabeled data—an assumption that may fail in low-resource settings.
Vision Foundation Models. Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, has become indispensable to modern AI development awais2025foundation . These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks. For example, CLIP-ViT radford2021learningtransferablevisualmodels trained with millions of image-text pairs shows an unprecedented zero-shot capability, robustness to distribution shifts, and serves as the encoders for various powerful generative models rombach2022high ; liu2024visual . Meanwhile, DINOv2 oquab2024dinov2learningrobustvisual pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features. Moreover, it is increasingly becoming a common sense that strategically fusing multiple VFMs can unlock synergistic gains mai2024lessonslearnedunifyingempirical ; tong2024eyes , leading to state-of-the-art performance across tasks from visual question answering to object detection.
@@ 3.2 A Comprehensive SSL Image Classification Benchmark
Despite recent advancements in SSL, most studies gan2024erasingbiasfinetuningfoundation ; chen2023softmatch ; zhang2021flexmatch continue to evaluate on classic datasets such as CIFAR-10/100 krizhevsky2009learning , STL-10 pmlr-v15-coates11a , and Food101 bossard2014food . However, these benchmarks exhibit two key limitations: Diminishing difficulty under VFMs. Following prior PEFT-on-VFM work jie2022convolutionalbypassesbettervision ; jia2022visual ; tu2023visual ; mai2024lessonslearnedunifyingempirical , we use linear probing as our primary performance indicator. As shown in Table 1, linear probing on frozen VFM backbones already delivers remarkable accuracy—even with limited labeled data. We also report results for various classification-head sizes in section B.4, and Narrow domain coverage. Because these benchmarks focus mainly on natural images, they offer only a narrow view of real-world VFM applications. To effectively evaluate SSL methods in the VFM era—focusing exclusively on semi-supervised image classification—we propose a new benchmark designed to capture the complexities of real-world applications across different domains and dataset sizes.
Dataset & Regime We choose six classification datasets from VTAB zhai2020largescalestudyrepresentationlearning , spanning remote sensing, medical imaging, synthetic data, and autonomous driving tasks. To test the robustness of SSL methods, we vary the number of labels per class and employ linear probing as our performance metric, selecting shot counts that keep each task sufficiently challenging for frozen VFMs. These settings impose substantial difficulty on frozen VFMs (Table 1), underscoring the importance of SSL for unlocking their full potential. A summary of the datasets appears in Table 2, with full details provided in Appendix A. Together, this benchmark offers a more diverse and comprehensive foundation-model-era evaluation of semi-supervised image classification.
@@ 3.3 Fair Hyperparameter Tuning for SSL
Dataset Task Domain $|\mathcal{L}|+|\mathcal{U}|$ Classes $|\mathcal{L}|/$Class
@@
However, ericsson2023betterpracticesdomainadaptation also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios. Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI vinh2009information , ARI halkidi2002cluster , V-Measure rosenberg2007v , FMI fowlkes1983method and BNM cui2020towards , and two from logits, RankMe garrido2023rankmeassessingdownstreamperformance and CHI calinski1974dendrite —for more robust tuning. Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{V}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{|\mathcal{V}|}$. Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all. We provide formal definitions and detailed procedures in section C.1 and the effectiveness of the proposed method in section B.7. By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@ 4 Systematic Evaluation of SSL with VFMs
Given the struggling performance of frozen VFMs in our benchmark (Table 1), we consider two straightforward strategies to enhance their effectiveness: (1) exploit the inherent generalizability of VFMs by fine-tuning only on labeled data; (2) employ SSL to leverage both labeled and unlabeled data. With the diverse benchmark and hyperparameter tuning protocol introduced in section 3, we aim to investigate whether existing SSL algorithms remain effective when adopting VFMs as their backbone.
@@
Can PEFT Come to the Rescue? When labeled downstream data are scarce, recent research has shown that parameter-efficient fine-tuning (PEFT)—which updates only a small subset of parameters or introduces a lightweight learnable module to frozen VFMs—often outperforms full fine-tuning on VFMs mai2024lessonslearnedunifyingempirical ; vpetlbench2024 ; hu2021loralowrankadaptationlarge ; chen2022adaptformeradaptingvisiontransformers ; jia2022visual ; tu2023visual . When considering SSL scenarios with limited labeled data, a natural question arises: is PEFT compatible with SSL? Although a recent study has made an initial attempt to apply PEFT in SSL gan2024erasingbiasfinetuningfoundation , a comprehensive analysis of its compatibility with different VFM backbones, PEFT methods, and SSL approaches is still lacking. We, therefore, assess two commonly used PEFT strategies: LoRA hu2021loralowrankadaptationlarge , which trains additive low-rank matrices for Transformer layers to approximate weight updates, and AdaptFormer chen2022adaptformeradaptingvisiontransformers , an adapter-based approach proven effective in computer vision. We follow the same protocol outlined in section 3 to tune the hyperparameters of these PEFT methods.
@@ 5.1 Self-training
[…]
and add these pseudo-labeled data to the labeled set =$\mathcal{L}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{|\mathcal{L}|}$ for supervised learning. Here, $\bm{\theta}$ represents the weights of the current classifier, and $\tau$ is the confidence threshold. If (a) high-confidence predictions are correct and (b) the updated model gradually increases its confidence in the remaining data, self-training can be as effective as fully supervised learning on the entire dataset with true labels.
@@ 5.2 Ensembling of multiple PEFT and VFMs
[…]
@@ 6 Experiment
Figure 3: Average SSL accuracy with full fine-tuning or PEFT across 12 settings shows that, with fair hyper-parameter tuning, fine-tuning on limited labels can outperform SSL; PEFT boosts SSL yet matches labeled-only performance, indicating minimal unlabeled-data benefit in current VFM-based SSL (see section B.3).
@@ 6.1 Performance Comparison
We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in subsection 6.1. Overall, V-PET achieve superior performance compared with others, as shown in the Average column. To further analyze the results, Figure 6 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings. We then compute the mean rank (number in the bracket) of each method and sort them accordingly. Although V-PET does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
@@
Table 3: Performance comparison of baselines, existing SSL approaches, and our proposed methods on six diverse datasets (12 settings). The best result within each PEFT and VFM is highlighted in bold. We report the full results in section B.2.
@@ 6.2 Scalability of Ensembling
Impact of Pseudo-Label Quality. When applying V-PET, one might wonder: if the performances of different VFMs vary significantly, can their pseudo-labels still be ensembled to boost downstream results? The answer is yes, we demonstrate the result in section B.6.
@@ B.2 SSL and Labeled-Only Baseline Details
As a completion of subsection 6.1, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in subsection B.2.
@@ B.5 Impact of Labeled-Only Initialization on SSL
Per discussion in section 4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT. To explore this, we conduct experiments on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10). We first train DINOv2 models on labeled data using LoRA and then fine-tune them with SoftMatch. These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in subsection 6.1.
@@ C.1 Hyperparameter Tuning Procedure
[…]
@@
To get the best performance model, we select the model with the lowest value in $A$. We put the validators we use and more detailed examples in section C.2.
@@ C.2 Validators
$CHI\coloneqq\frac{\left[\sum_{i=1}^{k}n_{i}||c_{i}-c||^{2}\right]/(k-1)}{\left[\sum_{i=1}^{k}\sum_{x\in C_{i}}||x-c_{i}||^{2}\right](n-k)}$

### v3

@@ 1 Introduction
Recently, pre-training on external labeled or unlabeled data has become the de facto standard across many machine learning application domains gui2024survey ; mai2024fine . For example, in computer vision, many recent algorithms are built upon vision foundation models (VFMs) such as CLIP radford2021learningtransferablevisualmodels and DINOv2 oquab2024dinov2learningrobustvisual . These models were pre-trained on massive datasets—millions, if not billions, of data points. As a result, they have demonstrated remarkable generalizability across a wide range of tasks, often requiring only slight fine-tuning or, in some cases, functioning effectively as a frozen backbone zhang2024vision .
@@ 2 Related Works
Transfer Learning & Self-Supervised Learning. Transfer learning zhuang2020comprehensivesurveytransferlearning ; tu2023holistic —and specifically PEFT (or PETL) for efficiently adapting foundation models han2024parameterefficientfinetuninglargemodels —has long leveraged pretrained models to boost downstream tasks. In our approach, we fine-tune VFMs via PEFT on labeled data, then use the adapted models to generate pseudo-labels for unlabeled examples, thus bridging transfer learning with semi-supervised methods. A related paradigm, self-supervised learning, also exploits unlabeled data and often serves as a baseline in SSL comparisons chen2020bigselfsupervisedmodelsstrong ; assran2022maskedsiamesenetworkslabelefficient , but it assumes abundant unlabeled data—an assumption that may fail in low-resource settings.
Vision Foundation Models. Vision Transformers (ViT) dosovitskiy2021imageworth16x16words pre-trained on massive amounts of data, has become indispensable to modern AI development awais2025foundation ; mai2025ava ; gu2025bioclip . These models, often referred to as vision foundation models (VFMs), have demonstrated superior performance on a wide range of tasks. For example, CLIP-ViT radford2021learningtransferablevisualmodels trained with millions of image-text pairs shows an unprecedented zero-shot capability, robustness to distribution shifts, and serves as the encoders for various powerful generative models rombach2022high ; liu2024visual . Meanwhile, DINOv2 oquab2024dinov2learningrobustvisual pre-trained with self-supervised objectives on extensive sets of well-curated images effectively captures fine-grained localization features. Moreover, it is increasingly becoming a common sense that strategically fusing multiple VFMs can unlock synergistic gains mai2024lessonslearnedunifyingempirical ; tong2024eyes , leading to state-of-the-art performance across tasks from visual question answering to object detection.
@@ 3.2 A Comprehensive SSL Image Classification Benchmark
Despite recent advancements in SSL, most studies gan2024erasingbiasfinetuningfoundation ; chen2023softmatch ; zhang2021flexmatch continue to evaluate on classic datasets such as CIFAR-10/100 krizhevsky2009learning , STL-10 pmlr-v15-coates11a , and Food101 bossard2014food . However, these benchmarks exhibit two key limitations: Diminishing difficulty under VFMs. Following prior PEFT-on-VFM work jie2022convolutionalbypassesbettervision ; jia2022visual ; tu2023visual ; mai2024lessonslearnedunifyingempirical ; xiao2025visual ; zhang2025ot , we use linear probing as our primary performance indicator. As shown in Table 1, linear probing on frozen VFM backbones already delivers remarkable accuracy—even with limited labeled data. We also report results for various classification-head sizes in section˜B.4, and Narrow domain coverage. Because these benchmarks focus mainly on natural images, they offer only a narrow view of real-world VFM applications. To effectively evaluate SSL methods in the VFM era—focusing exclusively on semi‐supervised image classification—we propose a new benchmark designed to capture the complexities of real‐world applications across different domains and dataset sizes.
Dataset & Regime We choose six classification datasets from VTAB zhai2020largescalestudyrepresentationlearning , spanning remote sensing, medical imaging, synthetic data, and autonomous driving tasks. To test the robustness of SSL methods, we vary the number of labels per class and employ linear probing as our performance metric, selecting shot counts that keep each task sufficiently challenging for frozen VFMs. These settings impose substantial difficulty on frozen VFMs (Table 1), underscoring the importance of SSL for unlocking their full potential. A summary of the datasets appears in Table 2, with full details provided in Appendix A. Together, this benchmark offers a more diverse and comprehensive foundation‐model‐era evaluation of semi‐supervised image classification.
@@ 3.3 Fair Hyperparameter Tuning for SSL
[…]
@@
However, ericsson2023betterpracticesdomainadaptation also revealed that no single criterion could reliably select suitable hyperparameters across all scenarios. Motivated by this finding, we propose integrating seven unsupervised criteria—five derived from features, AMI vinh2009information , ARI halkidi2002cluster , V-Measure rosenberg2007v , FMI fowlkes1983method and BNM cui2020towards , and two from logits, RankMe garrido2023rankmeassessingdownstreamperformance and CHI calinski1974dendrite —for more robust tuning. Concretely, for each hyperparameter configuration and its corresponding model, we compute all seven criteria using a held-out unlabeled validation set $\mathcal{V}\coloneqq\{{\bm{x}}^{v}_{i}\}_{i=1}^{|\mathcal{V}|}$. Next, we rank every hyperparameter configuration based on each criterion and choose the one achieving the lowest average rank across all. We provide formal definitions and detailed procedures in section˜C.1 and the effectiveness of the proposed method in section˜B.7. By eliminating reliance on held-out labeled validation sets, our method mitigates data leakage and promotes a more practical and fair tuning protocol for SSL.
@@ 4 Systematic Evaluation of SSL with VFMs
Given the struggling performance of frozen VFMs in our benchmark (Table 1), we consider two straightforward strategies to enhance their effectiveness: (1) exploit the inherent generalizability of VFMs by fine-tuning only on labeled data; (2) employ SSL to leverage both labeled and unlabeled data. With the diverse benchmark and hyperparameter tuning protocol introduced in section˜3, we aim to investigate whether existing SSL algorithms remain effective when adopting VFMs as their backbone.
@@
Can PEFT Come to the Rescue? When labeled downstream data are scarce, recent research has shown that parameter-efficient fine-tuning (PEFT)—which updates only a small subset of parameters or introduces a lightweight learnable module to frozen VFMs—often outperforms full fine-tuning on VFMs mai2024lessonslearnedunifyingempirical ; vpetlbench2024 ; hu2021loralowrankadaptationlarge ; chen2022adaptformeradaptingvisiontransformers ; jia2022visual ; tu2023visual . When considering SSL scenarios with limited labeled data, a natural question arises: is PEFT compatible with SSL? Although a recent study has made an initial attempt to apply PEFT in SSL gan2024erasingbiasfinetuningfoundation , a comprehensive analysis of its compatibility with different VFM backbones, PEFT methods, and SSL approaches is still lacking. We, therefore, assess two commonly used PEFT strategies: LoRA hu2021loralowrankadaptationlarge , which trains additive low-rank matrices for Transformer layers to approximate weight updates, and AdaptFormer chen2022adaptformeradaptingvisiontransformers , an adapter-based approach proven effective in computer vision. We follow the same protocol outlined in section˜3 to tune the hyperparameters of these PEFT methods.
@@ 5.1 Self-training
$\hat{y}^{u}_{i}\coloneqq\underset{c}{\operatorname{arg\,max}}\ p\left(c\big|{\bm{x}}^{u}_{i};f_{\bm{\theta}}\right)\hskip 20.0pt\text{ if }\ \ \underset{c}{\max}\ p(c\big|{\bm{x}}^{u}_{i};f_{\bm{\theta}})\geq\tau,$
and add these pseudo-labeled data to the labeled set $\mathcal{L}=\{({\bm{x}}^{l}_{i},y^{l}_{i})\}_{i=1}^{|\mathcal{L}|}$ for supervised learning. Here, $\bm{\theta}$ represents the weights of the current classifier, and $\tau$ is the confidence threshold. If (a) high-confidence predictions are correct and (b) the updated model gradually increases its confidence in the remaining data, self-training can be as effective as fully supervised learning on the entire dataset with true labels.
@@ 5.2 Ensembling of multiple PEFT and VFMs
$\mathcal{P}_{n,m}=\left\{\text{one\_hot}\bigl(\underset{c}{\operatorname{arg\,max}}f_{\tilde{\bm{\theta}}_{n,m}}(u)\bigr)\;\middle|\;\forall u\in\mathcal{U}\right\}.$
@@ 6 Experiment
Figure 3: Average SSL accuracy with full fine-tuning or PEFT across 12 settings shows that, with fair hyper-parameter tuning, fine-tuning on limited labels can outperform SSL; PEFT boosts SSL yet matches labeled-only performance, indicating minimal unlabeled-data benefit in current VFM-based SSL (see section˜B.3).
@@ 6.1 Performance Comparison
We summarize the performance comparison between ST, PET, V-PET and existing SSL methods across all the 12 settings in Table 3. Overall, V-PET achieve superior performance compared with others, as shown in the Average column. To further analyze the results, Figure 6 visualize the ranking frequency, where each matrix entry $(i,j)$ indicates how often method $i$ is ranked $j$-th across 12 settings. We then compute the mean rank (number in the bracket) of each method and sort them accordingly. Although V-PET does not claim first place in every single setting, it secures the top rank most frequently, showcasing the effectiveness of pseudo-label ensemble and establishing it as a simple yet powerful baseline in the VFM era.
@@
Table 3: Performance comparison of baselines, existing SSL approaches, and our proposed methods on six diverse datasets (12 settings). The best result within each PEFT and VFM is highlighted in bold. We report the full results in section˜B.2.
@@ 6.2 Scalability of Ensembling
Impact of Pseudo-Label Quality. When applying V-PET, one might wonder: if the performances of different VFMs vary significantly, can their pseudo-labels still be ensembled to boost downstream results? The answer is yes, we demonstrate the result in section˜B.6.
@@ 7 Conclusion
Acknowledgment
This research is supported by grants from the National Science Foundation (ICICLE: OAC-2112606). We appreciate the generous support of computational resources from the Ohio Supercomputer Center (OSC).
@@ B.2 SSL and Labeled-Only Baseline Details
As a completion of Table 3, we present the detailed results of the SSL baseline experiment with different SSL methods, including full fine-tuning, in Table 6.
@@ B.5 Impact of Labeled-Only Initialization on SSL
Per discussion in section˜4, a natural question is whether the performance of SSL methods can be improved by initializing from a model trained on labeled data using PEFT. To explore this, we conduct experiments on three datasets: DTD (N6), Diabetic Retinopathy (N80), and KITTI (N10). We first train DINOv2 models on labeled data using LoRA and then fine-tune them with SoftMatch. These datasets were chosen because their labeled-only performance exceeds that of SoftMatch by a large margin, as shown in Table 3.
@@ C.1 Hyperparameter Tuning Procedure
$f$ $\coloneqq g\circ h\in\mathbb{R}^{i}\rightarrow\mathbb{R}^{j}\quad$ $\text{(model)},$
$h$ $\in\mathbb{R}^{i}\rightarrow\mathbb{R}^{k}\quad$ $\text{(feature extractor)},$
$g$ $\in\mathbb{R}^{k}\rightarrow\mathbb{R}^{j}\quad$ $\text{(linear projector)}.$
@@
To get the best performance model, we select the model with the lowest value in $A$. We put the validators we use and more detailed examples in section˜C.2.
@@ C.2 Validators
[…]